### v2

# Enhancing Sum Capacity via Quantum and No-Signaling Cooperation Between Transmitters

Abstract
We consider a communication scenario over a discrete memoryless interference channel or multiple access channel without feedback, where transmitters exploit classical, quantum, or no-signaling cooperation. In this scenario, several previous works have shown that the sum capacities of channels involving pseudo-telepathy games can be enhanced by quantum or no-signaling cooperation. However, a full characterization of which channels admit such an improvement remains open. By focusing on the common characteristics of previously studied channels, we propose a broader class of channels for which quantum or no-signaling cooperation increases the sum capacity. Channels in this class are associated with a pseudo-telepathy game, with channel inputs specified as tuples of questions and answers from the game. In addition, when the channel inputs satisfy the winning condition of the game, the channel decomposes into parallel weakly symmetric sub-channels and is less noisy compared to the case when the inputs do not meet the winning condition.
I Introduction
Quantum information processing has attracted significant attention due to its potential to outperform classical information processing in certain tasks, thus exhibiting quantum advantage. Quantum advantage is not only of theoretical interest but also provides justification for the practical implementation of quantum information processing. Examples of quantum advantage arise in the contexts of algorithmic complexity [shorPolynomialTimeAlgorithmsPrime1997, groverFastQuantumMechanical1996, deutsch_rapid_1992], security [bennettQuantumCryptographyPublic2014, ekertQuantumCrypt1991, bennett_quantum_1992], privacy [namQuantumAdvantagePrivate2025], non-local games [brassardQuantumPseudoTelepathy2005a, cleveConsequencesLimitsNonlocal2004, brunnerBellNonlocality2014], distributed computing [cleve_substituting_1997, bar-yossef_exponential_2004, buhrmanNonlocalityCommunicationComplexity2010a], and communication over networks [leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, YunRaiBae2020, yunNonLocalQuantumAdvantages2023].
For quantum advantage in communication over networks, the works [leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, YunRaiBae2020, seshadriSeparationCorrelationAssistedSum2023, yunNonLocalQuantumAdvantages2023] considered scenarios where multiple transmitters cooperate based on pre-shared resources in transmitting classical messages over classical interference channels (IC) or multiple access channels (MAC). In this scenario, quantum advantage was demonstrated by showing that the sum capacity increases when the transmitters exploit quantum cooperation, as compared to classical cooperation, for specific channels involving pseudo-telepathy games [brassardQuantumPseudoTelepathy2005a, cleveConsequencesLimitsNonlocal2004, brunnerBellNonlocality2014]. Beyond quantum cooperation, no-signaling cooperation has also been considered to increase the sum capacity [quekQuantumSuperquantumEnhancements2017, YunRaiBae2020, yunNonLocalQuantumAdvantages2023, seshadriSeparationCorrelationAssistedSum2023]. While there have also been studies exploring more general scenarios involving no-signaling cooperation among both transmitters and receivers, as well as network channel models beyond MACs and ICs [fawziBroadcastChannelCoding2024, fawziMultipleAccessChannelCoding2024, notzelEntanglementEnabledCommunication2020], we focus on MACs and ICs with quantum and no-signaling cooperation between transmitters, as considered in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023]. Even in these simpler settings, it remains an open question under which classes of channels such an improvement can be achieved.
In the aforementioned settings, previous results in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023] were derived by adapting the characteristics of pseudo-telepathy games to the context of network communication. Specifically, a pseudo-telepathy game is a non-local game whose winning probability is strictly less than one with classical cooperation but equal to one with quantum or no-signaling cooperation. Examples of pseudo-telepathy games include the Clauser-Horne-Shimony-Holt (CHSH) game [clauserProposedExperimentTest1969], the magic square game [aravindBellsTheoremInequalities2002], and the multiparty parity game [brassardQuantumPseudoTelepathy2005a]. These games exhibit a hierarchy among three types of cooperation: quantum cooperation produces correlations unattainable by classical cooperation, and no-signaling cooperation generates correlations unattainable by quantum cooperation. Interpreted in the context of communication, quantum or no-signaling cooperation between transmitters can generate correlated channel inputs that are not possible through classical cooperation. Consequently, pseudo-telepathy games give rise to network channels in which interference between channel inputs can be mitigated by quantum or no-signaling cooperation, but not by classical cooperation.
The previously considered channels possessing the above properties are as follows. The works [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023] defined ICs and MACs where channel inputs are specified as tuples of questions and answers from a pseudo-telepathy game. Additionally, the channels studied in [leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017] relay the questions from transmitters to one or more receivers noiselessly when the channel inputs satisfy the winning condition of the game. More general channels have also been considered, where the transmitted questions are subject to random erasure [quekQuantumSuperquantumEnhancements2017, YunRaiBae2020], or to uniform random noise [yunNonLocalQuantumAdvantages2023], even when the channel inputs satisfy the winning condition. However, it has not yet been fully determined for which channels the sum capacity can be increased by quantum or no-signaling cooperation.
In this work, we propose a broader class of channels by focusing on the common characteristics shared by the previously studied channels: the interference is weaker when the channel inputs satisfy the winning condition of a pseudo-telepathy game than when they do not. Then, we show the enhancement of the sum capacity through quantum or no-signaling cooperation for such class of channels. Specifically, we propose a class of channels involving pseudo-telepathy games that become less noisy and decompose into parallel weakly symmetric sub-channels when the channel inputs satisfy the winning condition of the game. This class of channels includes the channels studied in prior works [leditzkyPlayingGamesMultiple2020, seshadriSeparationCorrelationAssistedSum2023, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025] as well as new ones not considered in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023]. To show the enhancement, we first derive an upper bound on the sum capacity of the proposed channels, which cannot be achieved by classical cooperation. Next, we show that the sum rate equal to this upper bound can be achieved if the transmitters exploit cooperation that wins the associated pseudo-telepathy game with probability one.
The rest of this paper is organized as follows. In Section II, we introduce the communication scenario with and without cooperation between transmitters, as well as preliminaries on non-local games and the hierarchy of cooperation. In Section III, we propose a class of channels involving non-local games and prove the enhancement of sum capacity via quantum or no-signaling cooperation. Section IV concludes the paper with a discussion of future work.
II System model & Preliminaries
In this work, we propose a class of ICs and MACs whose sum capacities can be enhanced through quantum or no-signaling cooperation between transmitters. Since all results for ICs can be directly adapted to those for MACs by treating multiple receivers as a single receiver (see Remark 1 for details), we primarily focus on ICs throughout this paper. To illustrate our results, we begin with a review of the conventional communication scenario over an IC, where all transmitters and receivers do not cooperate. Then, we describe an extended scenario that allows cooperation between transmitters, which is the focus of this work. Finally, we introduce non-local games, which serve as the key component underlying the proposed class of channels.
Notations: For $K\in\mathbb{N}$, let $K]:=\{1,\ldots,K\}$. For finite sets $\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{K}$, we denote $\mathcal{X}_{[K]}:=\prod_{i\in[K]}\mathcal{X}_{i}$, $\mathcal{X}_{i}^{n}:=\prod_{j\in[n]}\mathcal{X}_{i,j}$ where $\mathcal{X}_{i,j}=\mathcal{X}_{i}$ for all $j\in[n]$, and $\mathcal{X}_{[K]}^{n}:=\prod_{j\in[n]}\mathcal{X}_{[K],j}$ where $\mathcal{X}_{[K],j}:=\mathcal{X}_{[K]}$. A constant (e.g. $x_{[K]}^{n}$) or a random variable (e.g. $X_{[K]}^{n}$) with a given index is regarded as belonging to the corresponding set (e.g. $\mathcal{X}_{[K]}^{n}$).
II-A Communication over IC without transmitter cooperation
The conventional communication scenario between $K$ transmitter-receiver pairs, $K\geq 2$, over a stationary discrete memoryless interference channel (DM-IC) without feedback has been extensively studied [elgamalNetworkInformationTheory2011]. A DM-IC is characterized as a conditional probability $P_{Y_{[K]}|X_{[K]}}$ from $\mathcal{X}_{[K]}$ to $\mathcal{Y}_{[K]}$, where $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$ are finite sets denoting an input alphabet at the $i$-th transmitter side and an output alphabet at the $i$-th receiver side, respectively. This channel is stationary and memoryless in the sense that $n$-times use of the channel is given by $P_{Y_{[K]}^{n}|X_{[K]}^{n}}=P_{Y_{[K]}|X_{[K]}}^{\otimes n}$, i.e.,
$P_{Y_{[K]}^{n}|X_{[K]}^{n}}(y^{n}_{[K]}|x_{[K]}^{n})=\prod\limits_{j=1}^{n}P_{Y_{[K],j}|X_{[K],j}}(y_{[K],j}|x_{[K],j}).$ (1)
For each $i\in[K]$, the $i$-th transmitter has its own message $M_{i}\in\mathcal{M}_{i}$ and transmits a codeword $X^{n}_{i}$ corresponding to $M_{i}$ over the channel in $n$ time instances $P_{Y_{[K]}^{n}|X_{[K]}^{n}}$. Then, each of the $i$-th receiver decodes the noisy channel output $Y_{i}^{n}$ into $\hat{M}_{i}$ to estimate the message $M_{i}$.
In this case, general communication strategy over a DM-IC $P_{Y_{[K]}|X_{[K]}}$ that transmitters and receivers can use is characterized by a $(n,R_{[K]})$-code, which is defined as a collection of
  1. 1.
    Message sets $\mathcal{M}_{i}:=\left[2^{\lceil nR_{i}\rceil}\right]$ for all $i\in[n]$,
  2. 2.
    Encoding functions (encoders) $f_{i}:\mathcal{M}_{i}\rightarrow\mathcal{X}_{i}^{n}$ for all $i\in[n]$,
  3. 3.
    Decoding functions (decoders) $g_{i}:\mathcal{Y}_{i}^{n}\rightarrow\mathcal{M}_{i}\cup\{e\}$ for all $i\in[n]$, where $e$ denotes an error symbol.
Note that for a given $(n,R_{[K]})$-code, we have $X^{n}_{i}=f_{i}(M_{i})$ and $\hat{M}_{i}=g_{i}(Y^{n}_{i})$ for all $i\in[n]$. We say that a rate tuple $R_{[K]}$ is achievable (for $P_{Y_{[K]}|X_{[K]}}$) if there exists a sequence of $(n,R_{[K]})$-codes such that the average error probability converges to zero, i.e.,
$\lim\limits_{n\rightarrow\infty}\mathrm{Pr}(M_{[K]}\neq\hat{M}_{[K]})=0.$ (2)
The capacity region $\mathcal{C}(P_{Y_{[K]}|X_{[K]}})$ of a DM-IC $P_{Y_{[K]}|X_{[K]}}$ is defined as the closure of the set of all achievable rate tuples, and the sum capacity $\mathcal{C}_{s}(P_{Y_{[K]}|X_{[K]}})$ is defined as
$\mathcal{C}_{s}(P_{Y_{[K]}|X_{[K]}})\\ :=\max\left\{\sum_{i\in[K]}R_{i}:R_{[K]}\in\mathcal{C}(P_{Y_{[K]}|X_{[K]}})\right\}.$ (3)
To date, a single-letter characterization of the capacity region and the sum capacity of a DM-IC is not known in general, even for $K=2$. The best-known inner bound of the capacity region of a general DM-IC is the Han-Kobayashi inner bound [hanNewAchievableRate1981, chongHanKobayashiRegion2008], and this inner bound is tight for some class of DM-ICs.
II-B Communication over IC with transmitter cooperation
Now, we consider a scenario where transmitters cooperate using some pre-shared resources, which has been studied in [quekQuantumSuperquantumEnhancements2017, YunRaiBae2020, hawellekInterferenceChannelEntangled2025]. With such cooperation, the transmitters can transmit correlated codewords, and the correlation between them can be useful for mitigating interference, and hence, enlarges the capacity region.
Cooperation between the transmitters is characterized by a conditional probability that can be classified based on the type of pre-shared resource: classical, quantum, or super-quantum resource [brunnerBellNonlocality2014]. Cooperation that utilizes a classical, quantum, or super-quantum resource is termed classical, quantum, or no-signaling cooperation, respectively, and is defined as follows.
Definition 1
A conditional probability $P_{A_{[K]}|Q_{[K]}}$ is called
  1. 1.
    a classical cooperation if there exist a finite set $\mathcal{V}$, a probability distribution $P_{V}$ supported on $\mathcal{V}$, and conditional distributions $P_{A_{i}|Q_{i},V}$ for all $i\in[K]$ satisfying
    $P_{A_{[K]}|Q_{[K]}}(a_{[K]}|q_{[K]})\\ =\sum\limits_{v\in\mathcal{V}}P_{V}(v)\prod\limits_{i=1}^{K}P_{A_{i}|Q_{i},V}(a_{i}|q_{i},v).$ (4)
  2. 2.
    a quantum cooperation if there exist a quantum state $\rho$ and measurements (positive operator-valued measures) $\{\Pi^{i}_{a_{i}|q_{i}}\}_{a_{i}\in\mathcal{A}_{i}}$ for all $q_{i}\in\mathcal{Q}_{i}$ and $i\in[K]$ satisfying
    $P_{A_{[K]}|Q_{[K]}}(a_{[K]}|q_{[K]})=\mathrm{Tr}\left(\left(\otimes_{i=1}^{K}\Pi^{i}_{a_{i}|q_{i}}\right)\rho\right).$ (5)
  3. 3.
    a no-signaling cooperation if for all $i\in[K]$,
    $P_{A_{i}|Q_{[K]}}(a_{i}|q_{[K]})\\ =P_{A_{i}|Q_{i}}(a_{i}|q_{i}).$ (6)
    Simply, for all $i\in[K]$, $(Q_{1},\ldots,Q_{i-1},Q_{i+1},\ldots,Q_{K})$ and $A_{i}$ are conditionally independent given $Q_{i}$.
We denote by $\mathcal{E}^{\mathrm{C}},\mathcal{E}^{\mathrm{Q}}$, and $\mathcal{E}^{\mathrm{NS}}$ the sets of all classical, quantum, and no-signaling cooperations, respectively.
Note that a quantum cooperation is not a classical cooperation only if a quantum state $\rho$ in (5) is entangled. Also, there is a strict hierarchy among the above types of cooperations in the containment relation [brunnerBellNonlocality2014], i.e.,
$\mathcal{E}^{\mathrm{C}}\subsetneq\mathcal{E}^{\mathrm{Q}}\subsetneq\mathcal{E}^{\mathrm{NS}}.$ (7)
In the communication scenario allowing cooperation between transmitters, a general communication strategy that transmitters and receivers can use is characterized by an $(n,R_{[K]},r)$-code. The definition of a $(n,R_{[K]},r)$-code is the same as that of a $(n,R_{[K]})$-code described in Section II-A, except that the code contains cooperative encoders $P_{X^{n}_{[K]}|M_{[K]}}$ in $\mathcal{E}^{r}$ instead of encoding functions. Accordingly, we say that a rate tuple $R_{[K]}$ is $r$-achievable if there exists a sequence of $(n,R_{[K]},r)$-codes satisfying (2), the $r$-capacity region $\mathcal{C}^{r}$ is the closure of the set of all $r$-achievable rate tuples, and the $r$-sum capacity $\mathcal{C}_{s}^{r}$ is the maximum of $\sum_{i\in[K]}R_{i}$ over all $R_{[K]}\in\mathcal{C}^{r}$.
In the case $K=2$, previous works [quekQuantumSuperquantumEnhancements2017, hawellekInterferenceChannelEntangled2025, YunRaiBae2020] showed that for certain DM-ICs related to pseudo-telepathy games, utilizing quantum or no-signaling cooperation achieves a higher sum capacity than classical cooperation, i.e., $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{NS}}$.
II-C Non-local games
The usefulness of cooperation in a higher hierarchy can be explained in a non-local game scenario [brassardQuantumPseudoTelepathy2005a, cleveConsequencesLimitsNonlocal2004, brunnerBellNonlocality2014]. A promise-free $K$-party game $G$ is defined by a tuple of finite sets $(\mathcal{Q}_{[K]},\mathcal{A}_{[K]},\mathcal{W})$ where $\mathcal{W}\subseteq\mathcal{Q}_{[K]}\times\mathcal{A}_{[K]}$. For each $i\in[K]$, $\mathcal{Q}_{i}$ and $\mathcal{A}_{i}$ denote the sets of questions and answers for the $i$-th player, respectively, and $\mathcal{W}$ denotes the winning condition. [^1] A promise is a subset $P$ of the set of questions $\mathcal{Q}_{[K]}$. For a $K$-party game $G$ with promise $P$, a referee selects questions $Q_{[K]}$ only from $P$. In this paper, we only consider a promise-free game, i.e., $P=\mathcal{Q}_{[K]}$.
In a scenario where $K$ players play a game $G$, all the players discuss and agree on a cooperation strategy $P_{A_{[K]}|Q_{[K]}}$ before the game starts. After the game starts, each player enters an isolated room, i.e., they cannot communicate with each other from now on. A referee then samples a uniform random questions $Q_{[K]}\in\mathcal{Q}_{[K]}$ and gives $Q_{i}$ to the $i$-th player for all $i\in[n]$. Based on the pre-determined cooperation strategy and the questions $Q_{[K]}$, each of the $i$-th player gives an answer $A_{i}$ to the referee. The referee declares that the players win if $(Q_{[K]},A_{[K]})\in\mathcal{W}$.
For a given game $G$, the usefulness of a cooperation $P_{A_{[K]}|Q_{[K]}}$ can be measured by the winning probability in the game. Accordingly, the usefulness of the types of cooperations $\mathcal{E}^{r}$ can be measured by the maximum winning probability over $P_{A_{[K]}|Q_{[K]}}\in\mathcal{E}^{r}$ in the game, defined as
$p_{w}^{r}(G):=\max\limits_{P_{A_{[K]}|Q_{[K]}}\in\mathcal{E}^{r}}\mathrm{Pr}((Q_{[K]},A_{[K]})\in\mathcal{W})$ (8)
$=\max\limits_{P_{A_{[K]}|Q_{[K]}}\in\mathcal{E}^{r}}\sum\limits_{(q_{[K]},a_{[K]})\in\mathcal{W}}\frac{P_{A_{[K]}|Q_{[K]}}(a_{[K]}|q_{[K]})}{|\mathcal{Q}_{[K]}|}.$ (9)
If $p_{w}^{r}(G)=1$ and $p_{w}^{\mathrm{C}}(G)<1$ for some $r\in\{\mathrm{Q},\mathrm{NS}\}$, then we call a game $G$ an $r$-pseudo-telepathy game, and a cooperation $P_{A_{[K]}|Q_{[K]}}\in\mathcal{E}^{r}$ that achieves $p_{w}^{r}(G)=1$ a winning strategy.
There are games that show strict gaps between the maximum winning probabilities that can be achieved by different types of cooperations. One of the most representative examples is the Clauser–Horne–Shimony–Holt (CHSH) game [clauserProposedExperimentTest1969]. The CHSH game $G_{\mathrm{CHSH}}$ is a 2-party game such that
$\forall i\in\{1,2\},\;\mathcal{Q}_{i}=\mathcal{A}_{i}=\{0,1\},$ (10)
and
$\mathcal{W}=\{(q_{[2]},a_{[2]}):a_{1}+a_{2}\equiv q_{1}q_{2}\;(\mathrm{mod}\;2)\}.$ (11)
The CHSH game is a $\mathrm{NS}$-pseudo-telepathy game, and
$p_{w}^{\mathrm{C}}(G_{\mathrm{CHSH}})=0.75$ $<p_{w}^{\mathrm{Q}}(G_{\mathrm{CHSH}})\approx 0.85$ (12)
$<p_{w}^{\mathrm{NS}}(G_{\mathrm{CHSH}})=1.$ (13)
Therefore, the CHSH game translates the strict hierarchy in the maximum winning probabilities to the strict hierarchy in the containment relation in (7).
One example of a $\mathrm{Q}$-pseudo-telepathy game is the magic square game $G_{\mathrm{MS}}$ [aravindBellsTheoremInequalities2002]. It is a two-party game that consists of
$\forall i\in\{1,2\},\;\mathcal{Q}_{i}=\{1,2,3\},\;\mathcal{A}_{i}=\{0,1\}^{3},$ (14)
and
$\mathcal{W}=\left\{(q_{[2]},a_{[2]}):\begin{matrix}\sum_{j\in[3]}a_{1,j}\equiv 0\;(\mathrm{mod}\;2),\\ \sum_{j\in[3]}a_{2,j}\equiv 1\;(\mathrm{mod}\;2),\\ a_{1,q_{2}}=a_{2,q_{1}}\end{matrix}\right\}.$ (15)
The magic square game also demonstrates a separation between winning probabilities,
$p_{w}^{\mathrm{C}}(G_{\mathrm{MS}})=8/9<p_{w}^{\mathrm{Q}}(G_{\mathrm{MS}})=1.$ (16)
Beyond two-party $\mathrm{Q}$-pseudo-telepathy game, there are $K$-party $\mathrm{Q}$-pseudo-telepathy games for all $K\geq 3$, known as the (promise-free version of) $K$-party parity ($K$-$P$) game [brassardQuantumPseudoTelepathy2005a, seshadriSeparationCorrelationAssistedSum2023]. The $K$-$P$ game $G_{K\mathrm{-PP}}$ consists of
$\forall i\in[K],\;\mathcal{Q}_{i}=\mathcal{A}_{i}=\{0,1\},$ (17)
and
$\mathcal{W}=\left\{(q_{[K]},a_{[K]}):\sum\limits_{i\in[K]}q_{i}\equiv 1\;(\mathrm{mod}\;2)\right\}$ (18)
$\quad\cup\left\{(q_{[K]},a_{[K]}):\begin{matrix}\sum\limits_{i\in[K]}q_{i}\equiv 0\;(\mathrm{mod}\;2),\\ \sum\limits_{i\in[K]}a_{i}\equiv\frac{1}{2}\sum\limits_{i\in[K]}q_{i}\;(\mathrm{mod}\;2)\end{matrix}\right\}.$ (19)
For $G_{K\mathrm{-PP}}$ and $K\geq 3$, we have
$p_{w}^{\mathrm{C}}(G_{K\mathrm{-PP}})$ $=\frac{3}{4}+2^{-(\lceil K/2\rceil+1)}$ (20)
$<p_{w}^{\mathrm{Q}}(G_{K\mathrm{-PP}})=1.$ (21)
III Main results
Our main result is on showing the enhancement of sum capacity according to the hierarchy of cooperation, i.e., $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{NS}}$, for a certain class of DM-ICs and DM-MACs. The proposed class includes all the channels previously used to demonstrate such separation in previous work [leditzkyPlayingGamesMultiple2020, seshadriSeparationCorrelationAssistedSum2023, peregMultipleAccessChannelEntangled2025, hawellekInterferenceChannelEntangled2025], as well as many other channels beyond the channels considered in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023].
Remark 1
A DM-IC $P_{Y_{[K]}|X_{[K]}}$ can be identified as a DM-MAC $P_{\tilde{Y}|X_{[K]}}$ by treating multiple receivers as a single receiver, i.e., $\tilde{\mathcal{Y}}=\mathcal{Y}_{[K]}$ and the single receiver observes the channel output $\tilde{Y}=Y_{[K]}$. With this identification, all results and proofs for DM-ICs presented in the rest of the paper apply directly to those for the corresponding DM-MACs. Therefore, we provide a detailed presentation of the results only for DM-ICs.
III-A DM-ICs from non-local games
Inspired by $K$-party games and the previous works [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023], we propose a class of DM-ICs for which quantum or no-signaling cooperation can achieve strictly greater sum capacity than classical cooperation. The basic idea is to design a DM-IC $P_{Y_{[K]}|X_{[K]}}$ that is less noisy when channel inputs $x_{[K]}$ satisfy the winning condition of a game $G$ than when they do not satisfy. Specifically, the conditional entropy $H(Y_{[K]}|X_{[K]}=x_{[K]})$ is smaller when $x_{[K]}\in\mathcal{W}$ than when $x_{[K]}\notin\mathcal{W}$. Further, we set channel input $x_{i}$ to a pair of question and answer $(q_{i},a_{i})$ for $G$, and design $P_{Y_{[K]}|X_{[K]}}$ such that it decomposes into parallel sub-channels, i.e., $P_{Y_{[K]}|X_{[K]}}=\prod_{i\in[K]}P_{Y_{i}|Q_{i}}$, whenever $x_{[K]}\in\mathcal{W}$. Here, $a_{[K]}$ is discarded if the winning condition is satisfied. By requiring each sub-channel $P_{Y_{i}|Q_{i}}$ to be weakly symmetric [coverElementsInformationTheory2012, Sec. 7.2], we can show the enhancements of sum capacity for some of these channels. The details are as follows.
Definition 2
A conditional probability $P_{Y|Q}$ is weakly symmetric if
  1. 1.
    $\sum_{q\in\mathcal{Q}}P_{Y|Q}(y|q)$ is constant for all $y\in\mathcal{Y}$,
  2. 2.
    For any $q,q^{\prime}\in\mathcal{Q}$, there exists a permutation $\sigma$ on $\mathcal{Y}$ such that $P_{Y|Q}(\cdot|q^{\prime})=P_{Y|Q}(\sigma(\cdot)|q)$.
Lemma 1
If $P_{Y|Q}$ is weakly symmetric, then the point-to-point channel capacity of $P_{Y|Q}$ is given by
$\mathcal{C}(P_{Y|Q}):=\max\limits_{P_{Q}}I(Q;Y)=\log|\mathcal{Y}|-H(Y|Q=q),$ (22)
for all $q\in\mathcal{Q}$. Moreover, the maximum is achieved when $P_{Q}$ is the uniform distribution on $\mathcal{Q}$, and $P_{Y}$ becomes the uniform distribution on $\mathcal{Y}$ when $P_{Q}$ is the uniform distribution on $\mathcal{Q}$.
Definition 3
For a given $K$-party game $G=(\mathcal{Q}_{[K]},\mathcal{A}_{[K]},\mathcal{W})$, a $G$-game channel $P_{Y_{[K]}|X_{[K]}}^{G}$ is a DM-IC such that $\mathcal{X}_{i}=\mathcal{Q}_{i}\times\mathcal{A}_{i}$, $\mathcal{Y}_{[K]}$ is finite, and satisfies
  1. 1.
    For all $x_{[k]}\in\mathcal{W}$ and $\tilde{x}_{[K]}\notin\mathcal{W}$,
    $H(Y_{[K]}|X_{[K]}=x_{[K]})<H(Y_{[K]}|X_{[K]}=\tilde{x}_{[K]}).$ (23)
  2. 2.
    If $(q_{[K]},a_{[K]})\in\mathcal{W}$, then
    $P^{G}_{Y_{[K]}|Q_{[K]},A_{[K]}}(y_{[K]}|q_{[K]},a_{[K]})\\ =\prod\limits_{i\in[K]}P^{G}_{Y_{i}|Q_{i}}(y_{i}|q_{i}),$ (24)
    and $P^{G}_{Y_{i}|Q_{i}}$ is weakly symmetric for all $i\in[K]$.
For a given $G$-game channel $P_{Y_{[K]}|X_{[K]}}^{G}$, there exist $h_{1}^{w},\ldots,h_{K}^{w}\in\mathbb{R}$ such that for all $x_{[K]}\in\mathcal{W}$,
$H(Y_{[K]}|X_{[K]}=(q_{[K]},a_{[K]}))$ $=\sum_{i\in[K]}H(Y_{i}|Q_{i}=q_{i})$ (25)
$=\sum_{i\in[K]}h_{i}^{w},$ (26)
since (24) holds and $P^{G}_{Y_{i}|Q_{i}}$ is weakly symmetric for all $i\in[K]$. Hence, (23) is equivalent to
$\sum_{i\in[K]}h_{i}^{w}<H(Y_{[K]}|X_{[K]}=\tilde{x}_{[K]}).$ (27)
Example 1
There exist many $G$-game channels for a fixed $G$. As an example, consider a DM-IC $P^{G}_{Y_{[K]}|X_{[K]}}$ that splits into $K$ parallel channels, $P^{G}_{Y_{[K]}|X_{[K]}}=\prod_{i\in[K]}P^{G,w}_{Y_{i}|Q_{i}}$ when $x_{[K]}\in\mathcal{W}$ and $P^{G}_{Y_{[K]}|X_{[K]}}=\prod_{i\in[K]}P^{G,l}_{Y_{i}|Q_{i}}$ otherwise, where for $c\in\{w,l\}$,
$P^{G,c}_{Y_{i}|Q_{i}}(y_{i}|q_{i})=\eta_{c}\delta_{y_{i},q_{i}}+\frac{1-\eta_{c}}{|\mathcal{Q}_{i}|},$ (28)
with $0\leq\eta_{l}<\eta_{w}\leq 1$. These are instances of $G$-game channels that are not considered in the previous works [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023] (together with the identification of DM-ICs as DM-ICs mentioned in Remark 1). The channels considered in the previous works [leditzkyPlayingGamesMultiple2020, seshadriSeparationCorrelationAssistedSum2023, peregMultipleAccessChannelEntangled2025, hawellekInterferenceChannelEntangled2025] to show the enhancements of sum capacity are instances of a game channel of this form with $\eta_{w}=1$ and $\eta_{l}=0$.
Remark 2
Previous work [yunNonLocalQuantumAdvantages2023] considered a class of DM-MACs $P_{\tilde{Y}|X_{[K]}}$ determined by a game $G$, such that $\mathcal{X}_{i}=(\mathcal{Q}_{i},\mathcal{A}_{i})$, $\tilde{\mathcal{Y}}=\mathcal{Q}_{[K]}$, and
$P_{\tilde{Y}|X_{[K]}}(y_{[K]}|q_{[K]},a_{[K]})\\ =\begin{cases}\eta_{w}\prod\limits_{i\in[K]}\delta_{y_{i},q_{i}}+\frac{1-\eta_{w}}{\prod\limits_{i\in[K]}|\mathcal{Q}_{i}|}&\text{if }x_{[K]}\in\mathcal{W}\\ \eta_{l}\prod\limits_{i\in[K]}\delta_{y_{i},q_{i}}+\frac{1-\eta_{l}}{\prod\limits_{i\in[K]}|\mathcal{Q}_{i}|}&\text{otherwise},\end{cases}$ (29)
where $0\leq\eta_{l}<\eta_{w}\leq 1$. The sum capacity of this DM-MAC can be enhanced by quantum or no-signaling cooperation. Note that this DM-MAC can be identified as a DM-IC, where each of the $i$-th receiver observes $Y_{i}$ (cf. Remark 1). However, it is not clear whether the sum capacity of the associated DM-IC can be enhanced. This is because a sum rate $I(X_{[K]};Y_{[K]})$ is not achievable in general for a DM-IC. The property of our game channel that it splits into parallel channels when $x_{[K]}\in\mathcal{W}$ plays a key role in proving the enhancement (see Proposition 5 and its proof for details).
III-B Enhancing sum capacity
We show that quantum or no-signaling cooperation achieves higher sum capacity than classical cooperation for a $G$-game channel, when $G$ is a $\mathrm{Q}$- or $\mathrm{NS}$-pseudo-telepathy game.
Theorem 2
Let $G$ be a $K$-party game and $P_{Y_{[K]}|X_{[K]}}^{G}$ be a $G$-game channel. For $r\in\{\mathrm{Q},\mathrm{NS}\}$, if $G$ is an $r$-pseudo-telepathy game, then
$\mathcal{C}_{s}^{\mathrm{C}}(P_{Y_{[K]}|X_{[K]}}^{G})<\mathcal{C}_{s}^{r}(P_{Y_{[K]}|X_{[K]}}^{G}).$ (30)
As introduced in Section II-C, there are $Q$-pseudo-telepathy $K$-party games for all $K\geq 2$ (the magic square game for $K=2$, and the $K$-party parity game for $K\geq 3$). Thus, Theorem 2 implies that, for all $K\geq 2$, there is a DM-IC $P_{Y_{[K]}|X_{[K]}}$ which shows an advantage of quantum cooperation over classical cooperation.
In the remainder of this subsection, we prove Theorem 2. We begin by deriving an upper bound on $\mathcal{C}_{s}^{r}$ of a $G$-game channel (the converse part). This upper bound is strict if $r=\mathrm{C}$ and $G$ is a pseudo-telepathy game. Next, we prove that this upper bound is $r$-achievable when $G$ is an $r$-pseudo-telepathy game by exploiting the cooperation corresponding to a winning strategy for $G$ (the achievability part). Combining the converse and achievability proofs, we establish Theorem 2.
III-B1 Converse
We propose an upper bound on the $r$-sum capacity of a game channel as follows.
Proposition 3
For all $G$-game channels $P_{Y_{[K]}|X_{[K]}}^{r}$ and $r\in\{\mathrm{C,Q,NS}\}$, we have
$\mathcal{C}_{s}^{r}\left(P_{Y_{[K]}|X_{[K]}}^{G}\right)\leq\sum\limits_{i\in[K]}(\log|\mathcal{Y}_{i}|-h_{i}^{w}).$ (31)
Moreover, the above inequality is strict when $r=\mathrm{C}$ and $G$ is a pseudo-telepathy game.
The proof of the above proposition utilizes the fact that classical cooperation does not enlarge the capacity region.
Lemma 4
For all DM-ICs $P_{Y_{[K]}|X_{[K]}}$, we have that $\mathcal{C}(P_{Y_{[K]}|X_{[K]}})=\mathcal{C}^{\mathrm{C}}(P_{Y_{[K]}|X_{[K]}})$.
Proof:
Let $P_{X_{[K]}^{n}|M_{[K]}}\in\mathcal{E}^{\mathrm{C}}$ be a cooperative encoder that is utilized in a $(n,R_{[K]},\mathrm{C})$-code. Since the average error probability $\mathrm{Pr}(M_{[K]}\neq\hat{M}_{[K]})$ is linear in $P_{X_{[K]}^{n}|M_{[K]}}$, one of the extreme points of $\mathcal{E}^{\mathrm{C}}$ achieves the minimum error probability. An extreme point of $\mathcal{E}^{\mathrm{C}}$ is equivalent to the collection of deterministic encoding functions $f_{1},\ldots,f_{K}$ that can be performed by the transmitters without cooperation. Hence, there exists an $(n,R_{[K]})$-code whose average error probability is not greater than that of the given $(n,R_{[K]},\mathrm{C})$-code. ∎
Now, we prove Proposition 3.
Proof:
For any sequence of $(n,R_{[K]},r)$-codes or $(n,R_{[K]})$-codes that achieves $R_{[K]}$, we have
$n$ $\sum\limits_{i\in[K]}R_{i}=H(M_{[K]})$ (32)
$\leq I(M_{[K]};Y_{[K]}^{n})+n\epsilon_{n}$ (33)
$\leq I(X_{[K]}^{n};Y_{[K]}^{n})+n\epsilon_{n}$ (34)
$\leq\sum\limits_{j\in[n]}I(X_{[K],j};Y_{[K],j})+n\epsilon_{n}$ (35)
$=\sum\limits_{j\in[n]}(H(Y_{[K],j})-H(Y_{[K],j}|X_{[K],j}))+n\epsilon_{n}$ (36)
$\leq n\sum\limits_{i\in[K]}\log|\mathcal{Y}_{i}|-\sum\limits_{j=1}^{n}H(Y_{[K],j}|X_{[K],j})+n\epsilon_{n},$ (37)
where (33) follows from Fano’s inequality with $\epsilon_{n}\rightarrow 0$ as $n\rightarrow\infty$. Also, the data-processing inequality implies (34), and (35) holds since the channel is memoryless. The conditional entropy terms in (37) can be bounded as follows:
$H(Y_{[K],j}|X_{[K],j})$
$=\sum\limits_{x_{[K],j}\in\mathcal{W}}P_{X_{[K],j}}(x_{[K],j})H(Y_{[K],j}|X_{[K],j}=x_{[K],j})$ (38)
$\quad+\sum\limits_{x_{[K],j}\not\in\mathcal{W}}P_{X_{[K],j}}(x_{[K],j})H(Y_{[K],j}|X_{[K],j}=x_{[K],j})$
$\geq P_{X_{[K],j}}(\mathcal{W})\sum\limits_{i\in[K]}h_{i}^{w}+(1-P_{X_{[K],j}}(\mathcal{W}))\sum\limits_{i\in[K]}h_{i}^{w}$ (39)
$=\sum\limits_{i\in[K]}h_{i}^{w},$ (40)
where (39) follows from (26) and (27). Thus, we have (31).
Now, suppose $r=\mathrm{C}$ and $G$ is a pseudo-telepathy game, i.e., there is no classical winning strategy. By Lemma 4, it is enough to consider a scenario that the transmitters do not cooperate, i.e., $X_{i,j}=(Q_{i,j},A_{i,j})$ and $X_{i^{\prime},j}=(Q_{i^{\prime},j},A_{i^{\prime},j})$ are independent if $i\neq i^{\prime}$. Hence, we have $P_{A_{[K],j}|Q_{[K],j}}\in\mathcal{E}^{\mathrm{C}}$. Since $G$ is a pseudo-telepathy game, we have $P_{X_{[K],j}}(\mathcal{W})<1$. Combining with (27) and (39), we can conclude that $H(Y_{[K],j}|X_{[K],j})>\sum_{i\in[K]}h_{i}^{w}$. ∎
Remark 3
If the inequality in (31) is strict when $r=\mathrm{Q}$ and $G$ is a $\mathrm{NS}$-pseudo-telepathy game with $p_{w}^{\mathrm{Q}}(G)<1$, then we can show $\mathcal{C}_{s}^{\mathrm{Q}}<\mathcal{C}_{s}^{\mathrm{NS}}$ for the corresponding $G$-game channel. However, it is challenging to prove whether the inequality in (31) is strict in such a case, based on the similar approach in the last part of the above proof. This is because, it is not clear whether $P_{Q_{[K]},A_{[K]}|M_{[K]}}\in\mathcal{E}^{\mathrm{Q}}$ implies $P_{A_{[K]}|Q_{[K]}}\in\mathcal{E}^{\mathrm{Q}}$ or not.
Figure 1: A schematic diagram of the proposed achievability scheme over a single use $(n=1)$ of a game channel $P^{G}_{Y_{[K]}|X_{[K]}}$, which exploits a winning strategy $P_{A_{[K]}|Q_{[K]}}$ for a pseudo-telepathy game $G$. For $n\geq 2$, the part in the dash-dot line is copied $n$ times, i.e., $P_{Y_{[K]}|Q_{[K]}}$ is replaced by $P_{Y_{[K]}^{n}|Q_{[K]}^{n}}=P_{Y_{[K]}|Q_{[K]}}^{\otimes n}$.
III-B2 Achievability
We show that the upper bound on the $r$-sum capacity in Proposition 3 is achievable if a game $G$ is an $r$-pseudo-telepathy game. The main idea of the proof is to design an $(n,R_{[K]},r)$-code that incorporates cooperative encoding based on a winning strategy of a game $G$. Using this code, $P^{G}_{Y_{[K]}|X_{[K]}}$ splits into parallel channels, i.e., $P^{G}_{Y_{[K]}|X_{[K]}}=\prod_{i\in[K]}P^{G}_{Y_{i}|Q_{i}}$, and each $i$-th message can be reliably communicated at a rate equal to the capacity of the $i$-th sub-channel $P^{G}_{Y_{i}|Q_{i}}$.
Proposition 5
If $G$ is an $r$-pseudo-telepathy game, then
$\mathcal{C}_{s}^{r}(P_{Y_{[K]}|X_{[K]}}^{G})=\sum\limits_{i\in[K]}(\log|\mathcal{Y}_{i}|-h_{i}^{w}).$ (41)
Proof:
We consider a $(n,R_{[K]},r)$-code depicted in Fig. 1. Each of the transmitters first encodes their message with local encoding functions $f_{i}:\mathcal{M}_{i}\rightarrow\mathcal{Q}_{i}^{n}$. Then, for the $j$-th channel, the transmitters exploit the winning strategy $P_{A_{[K]}|Q_{[K]}}\in\mathcal{E}^{r}$ of the game $G$ with $Q_{[K],j}$ to get $A_{[K],j}$, and transmit $X_{[K],j}=(Q_{[K],j},A_{[K],j})$ to the channel. Note that the cooperative encoding $P_{X^{n}_{[K]}|M_{[K]}}$ constructed in the above is in $\mathcal{E}^{r}$. Since $\mathrm{Pr}(X_{[K],j}\in\mathcal{W})=1$, each of the $j$-th channel becomes $K$ parallel channels from $\mathcal{Q}_{[K]}$ to $\mathcal{Y}_{[K]}$, i.e., $P_{Y_{[K],j}|X_{[K],j}}=\prod_{i\in[K]}P_{Y_{i,j}|Q_{i,j}}$. Here, $P_{Y_{[K],j}|Q_{[K],j}}$ does not depend on $j\in[n]$. Hence, the overall communication scenario is equivalent to the communication over the $n$-times use of parallel channels $\prod_{i\in[K]}P_{Y_{i}|Q_{i}}^{\otimes n}$ without cooperation between transmitters. Thus, for each $i\in[K]$, the $i$-th transmitter-receiver pair can reliably communicate the message $M_{i}$ with the rate equal to the point-to-point channel capacity $\mathcal{C}(P_{Y_{i}|Q_{i}})$ of the $i$-th sub-channel $P_{Y_{i}|Q_{i}}$. Finally, Lemma 1 implies that the sum rate in (41) is achievable. ∎
IV Conclusion & Discussions
We proposed a class of DM-ICs and DM-MACs that involve non-local games: the channel decomposes into parallel weakly symmetric sub-channels and becomes less noisy when the channel inputs satisfy the winning condition of the game. In cases where the underlying non-local game is a quantum or no-signaling pseudo-telepathy game, quantum or no-signaling cooperation between the transmitters enhances the sum capacity, respectively.
Let us also address interesting future directions. First, it would be worth quantifying to what extent the sum capacity can be increased by quantum or no-signaling cooperation. For this purpose, it is necessary to quantify the sum capacity without cooperation. As one approach, the generalized Blahut-Arimoto algorithm [rezaeianComputationTotalCapacity2004], proposed for computing the classical sum capacity of DM-MACs (which serves as an upper bound on the sum capacities of DM-ICs), could be employed. However, although this algorithm can evaluate the sum capacities of certain DM-MACs, it generally provides only lower bounds on the sum capacities [buhlerNoteCapacityComputation2011]. Subsequently, [seshadriSeparationCorrelationAssistedSum2023] proposed an algorithm that computes the sum capacities of two-transmitter DM-MACs within a given additive precision in quasi-polynomial time. Consequently, it would also be interesting to investigate efficient algorithms for computing the sum capacity in multi-transmitter scenarios.
Second, it would be desirable to demonstrate sum capacity enhancements for channels relevant to current practical communication systems. A promising direction for future work would be to determine whether quantum cooperation can enhance the sum capacity of practical channels, such as Gaussian ICs or MACs. We conjecture that the sum capacity of a Gaussian channel can be enhanced when the structure of channel gains between transmitters and receivers resembles that of a quantum pseudo-telepathy game.
Another research direction would be to derive a single-letter characterization of the capacity region of general DM-ICs or DM-MACs with cooperation between transmitters. In this context, [peregMultipleAccessChannelEntangled2025] derived a multi-letter characterization of the capacity region of a DM-MAC, and [hawellekInterferenceChannelEntangled2025] proposed multi-letter inner and outer bounds for the capacity region of a DM-IC. Compared to single-letter characterizations, multi-letter characterizations have the limitation that it is not clear how to compute them [elgamalNetworkInformationTheory2011]. A single-letter characterization would help characterize the set of all channels in which cooperation enlarges the capacity region.
Acknowledgments
This work was supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (No. RS-2024-00452156), in part by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2023-00229524, Quantum Certification and Its Applications to Quantum SW (QC&QSW)), and in part by the NRF grant funded by the Korea government (MSIT) (No. RS-2025-00561467).
[Uncaptioned image] Seung-Hyun Nam (Member, IEEE) received the B.S. and M.S. degrees in Electrical Engineering from Pohang University of Science and Technology (POSTECH), Pohang, South Korea, in 2018 and 2020, respectively, and Ph.D. degree in Electrical Engineering from Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea, in 2024. He is currently a post-doctoral researcher with Information & Electronics Research Institute, KAIST. His research interests include (both classical and quantum) information theory, statistical inference, differential privacy, and information theoretic security.
[Uncaptioned image] Hyun-Young Park (Graduate Student Member, IEEE) received the B.S. degree in Electrical Engineering from Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea. He is currently pursuing the Ph.D. degree in the School of Electrical Engineering at Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea. His research interests include information theory, differential privacy, and quantum information theory.
[Uncaptioned image] Jiyoung Yun received the B.S. and M.S. degrees in Applied Mathematics from Hanyang University, and the Ph.D. degree in Electrical Engineering from the Korea Advanced Institute of Science and Technology (KAIST). She is currently a postdoctoral researcher at KAIST. Her research interests include quantum information theory, network quantum information theory, classical and quantum channel capacities, and entanglement-assisted communication.
[Uncaptioned image] Ashutosh Rai (Member, IEEE) did his Ph.D. work [2009-2013] in theoretical physics from the S. N. Bose National Center for Basic Sciences (University of Calcutta, Kolkata, India). He has worked as a researcher at the University of Latvia (Latvia), International Institute of Physics (Brazil), and Slovak Academy of Sciences (Slovakia). He is currently working at the School of Electrical Engineering, Korea Advanced Institute of Science and Technology (KAIST). His research interest cover topics in foundation of quantum mechanics and quantum information theory.
[Uncaptioned image] Si-Hyeon Lee (Senior Member, IEEE) received the B.S. (summa cum laude) and Ph.D. degrees in electrical engineering from the Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea, in 2007 and 2013, respectively. She is currently an Associate Professor with the School of Electrical Engineering, KAIST. She was a Postdoctoral Fellow with the Department of Electrical and Computer Engineering, University of Toronto, Toronto, Canada, from 2014 to 2016, and an Assistant Professor with the Department of Electrical Engineering, Pohang University of Science and Technology (POSTECH), Pohang, South Korea, from 2017 to 2020. Her research interests include information theory, wireless communications, statistical inference, and machine learning. She was an IEEE Information Theory Society Distinguished Lecturer (2024-2025) and a TPC Co-Chair of IEEE Information Theory Workshop 2024. She is currently an Associate Editor for IEEE Transactions on Information Theory, an Associate Editor for IEEE Transactions on Communications, and a Guest Editor for IEEE Journal on Selected Areas in Communications (Special issue on Secure Communication, Sensing, and Computation in Future Intelligent Wireless Networks).
[Uncaptioned image] Joonwoo Bae (Member, IEEE) obtained a Ph.D. in Theoretical Physics from Universitat de Barcelona & ICFO-Institute of Photonic Sciences, Barcelona in 2007. He has worked at the Korea Institute for Advanced Study (KIAS), Centre for Quantum Technologies (CQT) in Singapore, the ICFO, Freiburg Institute for Advanced Studies (FRIAS) as a Junior Fellow, and Hanyang University. He is currently with the School of Electrical Engineering, Korea Advanced Institute of Science and Technology (KAIST). His research interests include secure quantum communication, entanglement applications, open quantum systems, quantum foundations, and their practical applications.